# Haldane model on the Sierpiński gasket

### v2

@@ Abstract
We investigate the topological phases of the Haldane model on the Sierpiński gasket. As a consequence of the fractal geometry, multiple fractal gaps arise. Additionally, a flat band appears, and due to a complex next-nearest neighbour hopping, this band splits and multiple topological flux-induced gaps emerge. Owing to the fractal nature of the model, conventional momentum-space topological invariants cannot be used. Therefore, we characterise the system’s topology in terms of a real-space Chern number. In addition, we verify the robustness of the topological states to disorder. Finally, we present phase diagrams for both a fractal gap and a flux-induced gap. Previous work on a similar system claims that fractality “squeezes” the well-known Haldane phase diagram. However, this result arises because a doubled system was considered with two Sierpiński gaskets glued together. We consider only a single copy of the Sierpiński gasket, keeping global self-similarity. In contrast with these previous results, we find intricate and complex patterns in the phase diagram of this single fractal. Our work shows that the fractality of the model greatly influences the phase space of these structures, and can drive topological phases in the multitude of fractal and flux-induced gaps, providing a richer platform than a conventional integer dimensional geometry.
@@ I Introduction
For a long time, topological phases of matter have been at the forefront of condensed-matter research. This is because of their distinct and exotic properties, such as metallic boundary modes in an otherwise insulating bulk [1, 2]. The first type of topological state, the quantum Hall effect, was discovered by von Klitzing in 1980 by applying a magnetic field perpendicular to a 2D electron gas and performing transport measurements at low temperatures [3]. In 1988, Haldane showed that the requirement of a strong magnetic field to observe the quantum Hall effect in two-dimensional (2D) systems could be relaxed to a requirement of broken time-reversal (TR) symmetry [4]. In 2005, this idea was generalised by Kane and Mele with the introduction of spin-orbit coupling (SOC) in graphene, leading to the quantum spin Hall effect. In their seminal work, they showed that SOC breaks TR symmetry for a single spin, but that this symmetry gets restored by the other spin. Theoretically, this opens a topological gap [5, 6]. Although they overestimated the strength of this effect, these results kick-started the field of topological insulators. Since then, topological phases have been extensively studied theoretically and experimentally in photonic [7, 8, 9], acoustic [10, 11], and electronic systems [5, 12, 13]. These systems fit in the ‘tenfold way’ classification, which predicts the existence of topological phases based solely on the spectral symmetries and dimension of a system [14].
More recently, this framework has been expanded to describe non-Hermitian systems [15, 16, 17, 18], non-equilibrium systems [19, 20, 21, 22], quasicrystals [23, 24, 25, 26] and crystalline topology [27, 28, 29, 30], to cite only a few. A limitation of the ‘tenfold way’ is that it only accounts for integer dimensions. However, it is known that fractals can have a non-integer Hausdorff dimension $d_{H}$ [31], and host a new type of symmetry: self-similarity, which means that a subset of the system is similar to the entire system. In its most extreme case, self-similarity becomes scale invariance, when the subset and the entire system are indistinguishable.
Some of the most commonly considered fractals are the Sierpiński carpet and the Sierpiński gasket [32, 33, 34, 35, 36, 37, 38]. In Fig. 1, the Sierpiński gasket is shown up to its fourth generation.
@@
In 2015, Shang et al. synthesised the first molecular realisation of this fractal [33]. A few years later, Kempkes et al. presented the first electronic quantum fractal using artificial atoms and showed that the local density of states followed the Hausdorff dimension of this fractal [35]. Later, the dynamics in a photonic Sierpiński gasket was studied and it was found that the diffusion exponent became the Hausdorff dimension after the photons met the first void of the fractal [39]. Very recently, a naturally occurring Sierpiński gasket-shaped enzyme was observed [40].
In the last few years, these fractals and their possibly non-integer dimensionality have been combined with the study of topology, as an expansion of the ‘tenfold way’. On the theoretical side, the (integer) quantum Hall effect on the Sierpiński gasket has been investigated by M. Brezezińska et al. in 2018 [41] and further expanded upon in Refs. [42, 43]. In 2020, the first discussion of the fractional quantum Hall effect was presented and the existence of anyons in these fractals was predicted [44]. Not much later, higher-order topology and superconductivity in a Sierpiński gasket were theoretically investigated [45]. Finally, in 2023, a Hubbard model was considered on a Sierpiński gasket and the effects of SOC and electron-electron interactions were revealed [46].
@@
In this work, we consider just one copy of the Sierpiński gasket, keeping the asymmetry and the global self-similarity intact, and show that the phase diagram actually consists of more intricate and complex patterns. We implement a Haldane-like tight-binding model on two different geometries, a 2D triangular flake and the fractal Sierpiński gasket, and show that the fractal system has two distinct types of band gaps, fractal gaps and flux-induced gaps. We investigate the behaviour of both types of gaps and find that they can host topological modes for a wide range of parameters. These topological states are revealed by calculating a topological invariant. We have adopted a real-space Chern number, as the usual momentum-space Chern numbers are not applicable due to the finite size of these structures. Using this topological invariant, we find topological phase diagrams for the two gaps and observe complex patterns. In addition, we verify the resilience of these topological states to disorder. Our results show the richness of topology combined with fractality
The outline of this paper is as follows. In Sec. II, the Haldane model is introduced, and the two geometries, a triangular flake and the Sierpiński gasket are discussed. Then, in Sec. III, we present their energy spectra, examine the spectral properties, and investigate the localisation of a few boundary states. In Sec. IV, we characterise the topological nature of this model. To this end, we discuss the difficulties of the usual momentum-space Chern markers on the considered geometries and introduce the real-space Chern number. Thereafter, we study the emergence of the different topological gaps in terms of the parameters of the model, present the phase diagrams of this model, and compare them with earlier results. Finally, we probe the robustness of some topological states by including an arbitrary on-site disorder in Sec. V.
@@ II The Fractal Haldane model
The Haldane model is a tight-binding model on a honeycomb lattice with a nearest-neighbour (NN) hopping, a staggered (Semenoff) mass, and a complex NNN hopping. The Semenoff mass breaks the sublattice symmetry, while the complex NNN hopping breaks time reversal (TR) symmetry, even in the absence of a net external magnetic field [4]. From the tenfold way, we know that this combination of symmetries allows for topological phases in 2D [14]. The corresponding Hamiltonian is
@@
As shown in Fig. 2, when |$\absolutevalue{M}<\absolutevalue{3\sqrt{3}\lambda\sin(\Phi)}$ |, the system is a topological insulator, which is characterised by a nonzero quantised Hall conductance.
@@
This means that if one introduces edges to the system, it will host gap-crossing modes, which are strongly localised on the edges. On the other hand, when |$\absolutevalue{M}>\absolutevalue{3\sqrt{3}\lambda\sin(\Phi)}$ |, the system is a normal insulator, with zero conductance everywhere. The transition between these phases goes through a conducting phase, in which the band gap is closed. This one-to-one relationship between the bulk properties and the behaviour at the edges is known as the bulk-boundary correspondence, and is captured by the TKNN formula [54],
@@
In contrast, we investigate two lattices with open boundary conditions (OBC); a triangular flake and the fractal Sierpiński gasket lattice. These systems are no longer translationally invariant, and therefore one cannot represent the Hamiltonian in $k$-space. Nevertheless, the bulk-boundary correspondence still holds, and one can predict the behaviour at the edges if one has a well-defined bulk.
@@ II.1 Geometries
Figure 3: OBC lattice realisations corresponding to (a) a triangular flake and (b) a second-generation Sierpiński gasket. Sublattices $A$ and $B$ are marked in red and blue, respectively. The NN hoppings are depicted as solid black lines and NNN hoppings are depicted as dashed lines with a colour based on the sublattice that they connect.
@@ III.1 Energy Spectra
In Fig. 4(a), the characteristic Dirac cone around the $K$-point of the bulk is depicted. In Fig. 4(b), the spectrum of a triangular flake is shown. At first glance, it might appear to be gapped because there seem to be no states around $E=0$. However, this is a finite-size effect and the gap vanishes in the thermodynamic limit. In the fractal system [Fig. 4(c)], we observe new spectral gaps between $E=0$ and $E=\pm 1t$, and around $E=\pm 2t$. These gaps are neither present in the spectrum of the bulk nor of the triangular flake; instead, they are a consequence of fractality and not a result of finite-size effects. We will refer to them as fractal gaps. These fractal gaps are marked in orange. Furthermore, we notice a central flat band in both OBC cases [Figs. 4(b) and (c)], which is caused by the bipartite nature of these lattices in the absence of a complex NNN hopping [57, 46, 53]. Additionally, in the case of the Sierpiński gasket, multiple flat bands are observed away from $E=0$. These are comprised of compact localised states, as discussed in Ref. [46]. In Figs. 4(d)-(f), analogous plots are made, but now including a complex NNN hopping and a staggered mass $\lambda=M=0.1t$, with complex NNN hopping phase $\Phi=\pi/2$. Under these conditions, the bulk becomes gapped, as shown in Fig. 4(d). Upon consideration of the OBC systems, three notable features become apparent.
@@
First, we see that the spectra of the triangular flake and the Sierpiński gasket [Figs. 4(e) and 4(f), respectively] are slightly asymmetric around $E=0$ (dashed grey horizontal lines). This is a consequence of a nonzero staggered mass and the unbalanced number of lattice sites in each sublattice [57]. However, there is a fundamental distinction to be made between the fractal and the triangular flake. In the thermodynamic limit of the latter, the ratio of the number of sites in each sublattice tends to one, resulting in a negligible difference. This is because of the difference in the number of sites scales with the total edge length, while the total number of sites scales with the area. On the other hand, for the Sierpiński gasket, the ratio between the number of sites in each sublattice tends to $1.5$, making this difference and the corresponding asymmetric energy spectrum not just an outer-edge effect, but an intrinsic property of this Sierpiński gasket construction.
@@
Although the fractal gaps are never closed and remain trivial for $\lambda=M=0.1t$, this does not mean that the fractal does not host topological states. Previous investigation of similar fractal systems showed that these new non-degenerate states can have a topological character [47, 48]. Some of the dispersed modes of the central band can form new “bulk bands” away from $E=0$. Therefore, the introduction of a Haldane flux will cause the emergence of new gaps in the spectrum of this fractal, which we will call flux-induced gaps (marked in grey). These flux-induced gaps could in principle host topological “in-gap” states and the spreading of the states in the energy spectrum shown in Fig. 4(f) suggests that this is the case.
The difference in behaviour between the triangular flake and the fractal system is striking. The triangular flake was not gapped for $\lambda=M=0$, but has a finite bulk gap populated by topological edge modes when $\lambda=M=0.1t$. On the contrary, the Sierpiński gasket had (multiple) flat band(s) and multiple fractal gaps for $\lambda=M=0$. For the parameters in Fig. 4(f), the fractal gaps have not been closed, but instead, the flat band has split and the resulting flux-induced gaps are populated by “in-gap” states. Only upon further increasing $\lambda$ the fractal gap will close and reopen too, now populated by “fractal gap”-crossing edge states.
@@ III.2 Edge States
Figure 5: The probability density $|\psi|^{2}$ of some states on a fifth-generation Sierpiński gasket, with a nonzero complex NNN hopping. Here, $\lambda=0.2t$ and $M=0.05t$. These states correspond to points lying in between the bands depicted in Fig. 4(f). Each state’s energy is depicted as a point of the corresponding colour in the energy spectrum. The coloured states show strong localisation on the different edges of the Sierpiński gasket. The grey states lack this edge localisation and are “bulk” states.
To corroborate our claims in the previous section, we briefly investigate the spatial profile of the states in the flux-induced gaps. In Fig. 5, a subset of these states is shown, distinguished by their strong localisation on the different edges of a fifth-generation Sierpiński gasket (coloured) or lack thereof (grey). As a consequence of the many voids of the Sierpiński gasket, the lattice contains more edges. Subsequently, edge modes (coloured) are observed to localise on different sets of edges, ranging from edges around the smallest voids to the outermost edges, and combinations of these.
IV Topological Characterisation
In this section, we will introduce the topological invariant that will be used to characterise the topology of the different (fractal or flux-induced) gaps of the Sierpiński gasket. Usual calculation methods of topological invariants assume translational invariance, which is not present in a fractal. Moreover, the notion of a bulk-boundary correspondence becomes ill-defined, as a Sierpiński gasket has no bulk, due to voids on each length scale. Therefore, we will consider the local real-space method proposed by Kitaev to calculate a topological invariant [58]. Previous studies have applied this invariant to characterise the quantum Hall effect in fractals. Here, good agreement between this invariant and the transverse Hall current (obtained by direct transport calculations) is observed [42, 43].

### v3

@@ Abstract
We investigate the topological phases of the Haldane model on the Sierpiński gasket. As a consequence of the fractal geometry, multiple fractal gaps arise. Additionally, a flat band appears, and due to a complex next-nearest-neighbor hopping, this band splits and multiple topological flux-induced gaps emerge. Owing to the fractal nature of the model, conventional momentum-space topological invariants cannot be used. Therefore, we characterize the system’s topology in terms of a real-space Chern number. In addition, we verify the robustness of the topological states to disorder. Finally, we present phase diagrams for both a fractal gap and a flux-induced gap. Previous work on a similar system claims that fractality “squeezes” the well-known Haldane phase diagram. However, this result arises because a doubled system was considered with two Sierpiński gaskets glued together. We consider only a single copy of the Sierpiński gasket, keeping global self-similarity. In contrast with these previous results, we find intricate and complex patterns in the phase diagram of this single fractal. Our work shows that the fractality of the model greatly influences the phase space of these structures, and can drive topological phases in the multitude of fractal and flux-induced gaps, providing a richer platform than a conventional integer dimensional geometry.
@@ I Introduction
For a long time, topological phases of matter have been at the forefront of condensed-matter research. This is because of their distinct and exotic properties, such as metallic boundary modes in an otherwise insulating bulk [1, 2]. The first type of topological state, the quantum Hall effect, was discovered by von Klitzing in 1980 by applying a magnetic field perpendicular to a two-dimensional (2D) electron gas and performing transport measurements at low temperatures [3]. In 1988, Haldane showed that the requirement of a strong magnetic field to observe the quantum Hall effect in 2D systems could be relaxed to a requirement of broken time-reversal (TR) symmetry [4]. In 2005, this idea was generalized by Kane and Mele with the introduction of spin-orbit coupling (SOC) in graphene, leading to the quantum spin Hall effect. In their seminal work, they showed that SOC breaks TR symmetry for a single spin, but that this symmetry gets restored by the other spin. Theoretically, this opens a topological gap in graphene [5, 6]. Although they overestimated the strength of this effect, these results kick-started the field of topological insulators. Since then, topological phases have been extensively studied theoretically and experimentally in photonic [7, 8, 9], acoustic [10, 11], and electronic systems [5, 12, 13]. These systems fit in the ‘tenfold way’ classification, which predicts the existence of topological phases based solely on the spectral symmetries and dimension of a system [14].
More recently, this framework has been expanded to describe non-Hermitian systems [15, 16, 17, 18], non-equilibrium systems [19, 20, 21, 22], quasicrystals [23, 24, 25, 26] and crystalline topology [27, 28, 29, 30], to cite only a few. A limitation of the ‘tenfold way’ is that it only accounts for integer dimensions. However, it is known that fractals can have a non-integer Hausdorff dimension $d_{H}$ [31], and host a new type of symmetry: self-similarity, which means that a subset of the system is similar to the entire system. In its most extreme case, self-similarity becomes scale invariance, when the subset and the entire system are indistinguishable. Some of the earliest works studying physical properties of fractals considered phase transitions [32, 33, 34, 35, 36].
Two of the most commonly considered fractals are the Sierpiński carpet and the Sierpiński gasket [32, 37, 38, 39, 40, 41, 42, 43, 44, 45]. In Fig. 1, the Sierpiński gasket is shown up to its fourth generation.
@@
In 2015, Shang et al. synthesized the first molecular realization of this fractal [38]. A few years later, Kempkes et al. presented the first electronic quantum fractal using artificial atoms and showed that the local density of states followed the Hausdorff dimension of this fractal [40]. Later, the dynamics in a photonic Sierpiński gasket was studied and it was found that the diffusion exponent became the Hausdorff dimension after the photons met the first void of the fractal [46]. Very recently, a naturally occurring Sierpiński gasket-shaped enzyme was observed [47].
In the last few years, these fractals and their possibly non-integer dimensionality have been combined with the study of topology, as an expansion of the ‘tenfold way’. On the theoretical side, the (integer) quantum Hall effect on the Sierpiński gasket has been investigated by M. Brezezińska et al. in 2018 [48] and further expanded upon in Refs. [49, 50, 51, 52]. In 2020, the first discussion of the fractional quantum Hall effect was presented and the existence of anyons in these fractals was predicted [53]. Not much later, superconductivity, higher-order topology and non-Hermitian topology were theoretically investigated in a Sierpiński gasket [54, 55, 56]. Finally, in 2023, a Hubbard model was considered on a Sierpiński gasket and the effects of SOC and electron-electron interactions were revealed [57].
@@
In this work, we consider just one copy of the Sierpiński gasket, keeping the asymmetry and the global self-similarity intact, and show that the phase diagram actually consists of more intricate and complex patterns. We implement a Haldane-like tight-binding model on two different geometries, a 2D triangular flake and the fractal Sierpiński gasket, and show that the fractal system has two distinct types of band gaps, fractal gaps and flux-induced gaps. We investigate the behavior of both types of gaps and find that they can host topological modes for a wide range of parameters. These topological states are revealed by calculating a topological invariant. We have adopted a real-space Chern number, as the usual momentum-space Chern numbers are not applicable due to the finite size of these structures. Using this topological invariant, we find topological phase diagrams for the two gaps and observe complex patterns. In addition, we verify the resilience of these topological states to disorder. Our results show the richness of topology combined with fractality.
The outline of this paper is as follows. In Sec. II, the Haldane model is introduced, and the two geometries, a triangular flake and the Sierpiński gasket are discussed. Then, in Sec. III, we present their energy spectra, examine the spectral properties, and investigate the localization of a few boundary states. In Sec. IV, we characterize the topological nature of this model. To this end, we discuss the difficulties of the usual momentum-space Chern markers on the considered geometries and introduce the real-space Chern number. Thereafter, we study the emergence of the different topological gaps in terms of the parameters of the model, present the phase diagrams of this model, and compare them with earlier results. Finally, we probe the robustness of some topological states by including an arbitrary on-site disorder in Sec. V.
@@ II The Fractal Haldane model
The Haldane model is a tight-binding model on a honeycomb lattice with a nearest-neighbor (NN) hopping, a staggered (Semenoff) mass, and a complex NNN hopping. The Semenoff mass breaks the sublattice symmetry, while the complex NNN hopping breaks time reversal (TR) symmetry, even in the absence of a net external magnetic field [4]. From the tenfold way, we know that this combination of symmetries allows for topological phases in 2D [14]. The corresponding Hamiltonian is
@@
As shown in Fig. 2, when |$\absolutevalue{M}<\absolutevalue{3\sqrt{3}\lambda\sin(\Phi)}$ |, the system is a topological insulator, which is characterized by a nonzero quantized Hall conductance.
@@
This means that if one introduces edges to the system, it will host gap-crossing modes, which are strongly localized on the edges. On the other hand, when |$\absolutevalue{M}>\absolutevalue{3\sqrt{3}\lambda\sin(\Phi)}$ |, the system is a normal insulator, with zero conductance everywhere. The transition between these phases goes through a conducting phase, in which the band gap is closed. This one-to-one relationship between the bulk properties and the behavior at the edges is known as the bulk-boundary correspondence, and is captured by the TKNN formula [65],
@@
In contrast, we investigate two lattices with open boundary conditions (OBC); a triangular flake and the fractal Sierpiński gasket lattice. These systems are no longer translationally invariant, and therefore one cannot represent the Hamiltonian in $k$-space. Nevertheless, the bulk-boundary correspondence still holds, and one can predict the behavior at the edges if one has a well-defined bulk.
@@ II.1 Geometries
Figure 3: OBC lattice realizations corresponding to (a) a triangular flake and (b) a second-generation Sierpiński gasket. Sublattices $A$ and $B$ are marked in red and blue, respectively. The NN hoppings are depicted as solid black lines and NNN hoppings are depicted as dashed lines with a color based on the sublattice that they connect.
@@ III.1 Energy Spectra
In Fig. 4(a), the characteristic Dirac cone around the $K$-point of the bulk is depicted. In Fig. 4(b), the spectrum of a triangular flake is shown. At first glance, it might appear to be gapped because there seem to be no states around $E=0$. However, calculations performed for generations $g=3,4,5$ indicate that this is a finite-size effect and the gap vanishes in the thermodynamic limit [68], see Appendix C. In the fractal system [Fig. 4(c)], we observe new spectral gaps between $E=0$ and $E=\pm 1t$, and around $E=\pm 2t$. These gaps are neither present in the spectrum of the bulk nor of the triangular flake. They are a consequence of fractality and remain robust in the thermodynamic limit (see Appendix C). We will refer to them as fractal gaps. These fractal gaps are marked in orange. Furthermore, we notice a central flat band in both OBC cases [Figs. 4(b) and (c)], which is caused by the bipartite nature of these lattices in the absence of a complex NNN hopping [69, 57, 64]. Additionally, in the case of the Sierpiński gasket, multiple flat bands are observed away from $E=0$. These are comprised of compact localized states, as discussed in Ref. [57]. In Figs. 4(d)-(f), analogous plots are made, but now including a complex NNN hopping and a staggered mass $\lambda=M=0.1t$, with complex NNN hopping phase $\Phi=\pi/2$. Under these conditions, the bulk becomes gapped, as shown in Fig. 4(d). Upon consideration of the OBC systems, three notable features become apparent.
@@
First, we see that the spectra of the triangular flake and the Sierpiński gasket [Figs. 4(e) and 4(f), respectively] are slightly asymmetric around $E=0$ (dashed gray horizontal lines). This is a consequence of a nonzero staggered mass and the unbalanced number of lattice sites in each sublattice [69]. However, there is a fundamental distinction to be made between the fractal and the triangular flake. In the thermodynamic limit of the latter, the ratio of the number of sites in each sublattice tends to one, resulting in a negligible difference. This is because of the difference in the number of sites scales with the total edge length, while the total number of sites scales with the area. On the other hand, for the Sierpiński gasket, the ratio between the number of sites in each sublattice tends to $1.5$, making this difference and the corresponding asymmetric energy spectrum not just an outer-edge effect, but an intrinsic property of this Sierpiński gasket construction.
@@
Although the fractal gaps are never closed and remain trivial for $\lambda=M=0.1t$, this does not mean that the fractal does not host topological states. Previous investigation of similar fractal systems showed that these new non-degenerate states can have a topological character [58, 59]. Some of the dispersed modes of the central band can form new “bulk bands” away from $E=0$. Therefore, the introduction of a Haldane flux will cause the emergence of new gaps in the spectrum of this fractal, which we will call flux-induced gaps (marked in gray). These flux-induced gaps could in principle host topological “in-gap” states and the spreading of the states in the energy spectrum shown in Fig. 4(f) suggests that this is the case.
The difference in behavior between the triangular flake and the fractal system is striking. The triangular flake was not gapped for $\lambda=M=0$, but has a finite bulk gap populated by topological edge modes when $\lambda=M=0.1t$. On the contrary, the Sierpiński gasket had (multiple) flat band(s) and multiple fractal gaps for $\lambda=M=0$. For the parameters in Fig. 4(f), the fractal gaps have not been closed, but instead, the flat band has split and the resulting flux-induced gaps are populated by “in-gap” states. Only upon further increasing $\lambda$ the fractal gap will close and reopen too, now populated by “fractal gap”-crossing edge states.
@@ III.2 Edge States
Figure 5: The probability density $|\psi|^{2}$ of some states on a fifth-generation Sierpiński gasket, with a nonzero complex NNN hopping. Here, $\lambda=0.2t$ and $M=0.05t$. These states correspond to points lying in between the bands depicted in Fig. 4(f). Each state’s energy is depicted as a point of the corresponding color in the energy spectrum. The colored states show strong localization on the different edges of the Sierpiński gasket. The gray states lack this edge localization and are “bulk” states.
To corroborate our claims in the previous section, we briefly investigate the spatial profile of the states in the flux-induced gaps. In Fig. 5, a subset of these states is shown, distinguished by their strong localization on the different edges of a fifth-generation Sierpiński gasket (colored) or lack thereof (gray). As a consequence of the many voids of the Sierpiński gasket, the lattice contains more edges. Subsequently, edge modes (colored) are observed to localize on different sets of edges, ranging from edges around the smallest voids to the outermost edges, and combinations of these.
IV Topological Characterization
In this section, we will introduce the topological invariant that will be used to characterize the topology of the different (fractal or flux-induced) gaps of the Sierpiński gasket. Usual calculation methods of topological invariants assume translational invariance, which is not present in a fractal. Moreover, the notion of a bulk-boundary correspondence becomes ill-defined, as a Sierpiński gasket has no bulk, due to voids on each length scale. Therefore, we will consider the local real-space method proposed by Kitaev to calculate a topological invariant [70]. Previous studies have applied this invariant to characterize the quantum Hall effect in fractals. Here, good agreement between this invariant and the transverse Hall current (obtained by direct transport calculations) is observed [49, 51].